# Some new cases of Zilber–Pink in $Y(1)^{3}$

### v1

@@ 5. Proving Theorems 3 and 4
for $1\leq k\leq\ell$, which form a Zariski open covering of $\mathfrak{C}_{\mathfrak{p}}$. Then the rigid spaces $\operatorname{red}^{-1}(Z_{\mathfrak{p},1}),\dotsc,\operatorname{red}^{-1}(Z_{\mathfrak{p},\ell})$ form an admissible covering of $C^{v\textrm{-}\mathrm{an}}$. Since $U_{v,k}\subset\operatorname{red}^{-1}(Z_{\mathfrak{p},k})$ and $U_{v,k}$ is disjoint from $\operatorname{red}^{-1}(Z_{\mathfrak{p},j})$ for $j\neq k$, it follows that $U_{v,1},\dotsc,U_{v,\ell}$ form an admissible covering of their union.
@@
Meanwhile, if $\mathfrak{p}_{v}\not\in|\mathfrak{V}|$, we shrink $U_{v,1},\dotsc,U_{v,\ell}$ to make them disjoint and ensure that they form an admissible covering. To that end, choose $r^{\prime}_{v}<\min\{r_{v,1},\dotsc,r_{v,\ell}\}$ and let $U_{v,k}^{\prime}$ denote the preimage of the closed disc $\bar{D}(0,r_{v}^{\prime},K_{v})$ inside $U_{v,k}$. Then the rigid spaces $U_{v,k}^{\prime}$ are quasi-compact, so $U_{v,j}^{\prime}\cap U_{v,k}^{\prime}$ are quasi-compact for each $j,k$. Hence, if $U_{v,j}^{\prime}\cap U_{v,k}^{\prime}\neq\emptyset$, the absolute value of the function $x^{v\textrm{-}\mathrm{an}}$ has a minimum value $R_{v,jk}$ on $U_{v,j}^{\prime}\cap U_{v,k}^{\prime}$. If $j\neq k$, then $s_{j}^{v\textrm{-}\mathrm{an}},s_{k}^{v\textrm{-}\mathrm{an}}\not\in U_{v,j}^{\prime}\cap U_{v,k}^{\prime}$, so $R_{v,jk}>0$. Let
@@
By construction, after replacing $U_{v,k}$ by the preimage of $D(0,r_{v},K_{v})$ in itself, the sets $U_{v,1},\dotsc,U_{v,\ell}$ are pairwise disjoint.
Since the sets $U_{v,k}^{\prime}$ are affinoid and there are finitely many of them, they form an admissible covering of their union. After the final replacement of $U_{v,k}$, we have $U_{v,k}\subset U_{v,k}^{\prime}$ and $U_{v,k}$ is disjoint from $U_{v,j}^{\prime}$ for $j\neq k$. Hence $U_{v,1},\dotsc,U_{v,\ell}$ form an admissible covering of their union.
Thus, for each $v\in\Sigma_{K,f}$, we have constructed $U_{v,1},\dotsc,U_{v,\ell}$ satisfying (ii), (iv) and (v), which are pairwise disjoint and form an admissible covering of their union. Furthermore, for all $\mathfrak{p}_{v}\in|\mathfrak{V}|$, we have used the sets $U_{v,k}$ from Lemma 19 without shrinking them, and these satisfy $r_{v}=1$. Thus (i) is satisfied.
To conclude, we note that every point of $(\mathbb{A}^{1})^{v\textrm{-}\mathrm{an}}$ has at most $\ell$ preimages under $x^{v\textrm{-}\mathrm{an}}$. On the other hand, since $U_{v,1},\dotsc,U_{v,\ell}$ are pairwise disjoint, every point of $D(0,r_{v},K_{v})$ has $\ell$ distinct preimages in $U_{v,1}\cup\dotsb\cup U_{v,\ell}$. Hence the union $U_{v,1}\cup\dotsb\cup U_{v,\ell}$ is equal to $(x^{v\textrm{-}\mathrm{an}})^{-1}(D(0,r_{v},K_{v}))$. ∎

### v2

@@ Appendix A Rigid analytic neighbourhoods
for $1\leq k\leq\ell$, which form a Zariski open covering of $\mathfrak{C}^{\prime}_{\mathfrak{p}}$. Then the rigid spaces $(\operatorname{red}_{v}^{\prime})^{-1}(Z_{\mathfrak{p},1}),\dotsc,(\operatorname{red}_{v}^{\prime})^{-1}(Z_{\mathfrak{p},\ell})$ form an admissible covering of $C^{v\textrm{-}\mathrm{an}}$. Since $U_{v,k}\subset(\operatorname{red}_{v}^{\prime})^{-1}(Z_{\mathfrak{p},k})$ and $U_{v,k}$ is disjoint from $(\operatorname{red}_{v}^{\prime})^{-1}(Z_{\mathfrak{p},j})$ for $j\neq k$, it follows that $U_{v,1},\dotsc,U_{v,\ell}$ form an admissible covering of their union.
@@
Next, suppose that $\mathfrak{p}_{v}\not\in|\mathfrak{V}|$. If necessary, we will shrink $U_{v,1},\dotsc,U_{v,\ell}$ to make them disjoint and an admissible covering. To that end, choose $r^{\prime}_{v}<\min\{r_{v,1},\dotsc,r_{v,\ell}\}$ and let $U_{v,k}^{\prime}$ denote the preimage of the closed disc $\bar{D}(0,r_{v}^{\prime},K_{v})$ inside $U_{v,k}$. Then the rigid spaces $U_{v,k}^{\prime}$ are quasi-compact, so $U_{v,j}^{\prime}\cap U_{v,k}^{\prime}$ are quasi-compact for each $j,k$. Hence, if $U_{v,j}^{\prime}\cap U_{v,k}^{\prime}\neq\emptyset$, the absolute value of the function $x^{v\textrm{-}\mathrm{an}}$ has a minimum value $R_{v,jk}$ on $U_{v,j}^{\prime}\cap U_{v,k}^{\prime}$. If $j\neq k$, then $s_{j}^{v\textrm{-}\mathrm{an}},s_{k}^{v\textrm{-}\mathrm{an}}\not\in U_{v,j}^{\prime}\cap U_{v,k}^{\prime}$, so $R_{v,jk}>0$. Let
@@
By construction, after replacing $U_{v,k}$ by the preimage of $D(0,r_{v},K_{v})$, the sets $U_{v,1},\dotsc,U_{v,\ell}$ are pairwise disjoint.
Since the sets $U_{v,k}^{\prime}$ are affinoid and there are finitely many of them, they form an admissible covering of their union. After the final replacement of $U_{v,k}$, we have $U_{v,k}\subset U_{v,k}^{\prime}$ and $U_{v,k}$ is disjoint from $U_{v,j}$ for $j\neq k$. Hence $U_{v,1},\dotsc,U_{v,\ell}$ form an admissible covering of their union.
To conclude, we note that every point of $(\mathbb{A}^{1})^{v\textrm{-}\mathrm{an}}$ has at most $\ell$ preimages under $x^{v\textrm{-}\mathrm{an}}$. On the other hand, since $U_{v,1},\dotsc,U_{v,\ell}$ are pairwise disjoint, every point of $D(0,r_{v},K_{v})$ has $\ell$ distinct preimages in $U_{v,1}\cup\dotsb\cup U_{v,\ell}$. Hence the union $U_{v,1}\cup\dotsb\cup U_{v,\ell}$ is equal to $(x^{v\textrm{-}\mathrm{an}})^{-1}(D(0,r_{v},K_{v}))$.
Since the open subpaces $U_{v,k}$ constructed in this proof are always contained in the corresponding open subspaces $U_{v,k}$ obtained from Proposition A.2, they are simple. ∎